# Robustness-Aware 3D Object Detection in Autonomous Driving: A Review and Outlook

### v1

@@ Abstract
In the realm of modern autonomous driving, the perception system is indispensable for accurately assessing the state of the surrounding environment, thereby enabling informed prediction and planning. Key to this system is 3D object detection methods, that utilize vehicle-mounted sensors such as LiDAR and cameras to identify the size, category, and location of nearby objects. Despite the surge in 3D object detection methods aimed at enhancing detection precision and efficiency, there is a gap in the literature that systematically examines their resilience against environmental variations, noise, and weather changes. This study emphasizes the importance of robustness, alongside accuracy and latency, in evaluating perception systems under practical scenarios. Our work presents an extensive survey of camera-based, LiDAR-based, and multimodal 3D object detection algorithms, thoroughly evaluating their trade-off between accuracy, latency, and robustness, particularly on datasets like KITTI-C and nuScenes-C to ensure fair comparisons. Among these, multimodal 3D detection approaches exhibit superior robustness and a novel taxonomy is introduced to reorganize its literature for enhanced clarity. This survey aims to offer a more practical perspective on the current capabilities and constraints of 3D object detection algorithms in real-world applications, thus steering future research towards robustness-centric advancements.
@@
Autonomous Driving, 3D Object Detection, Point cloudsI Introduction
@@
AUTONOMOUS driving systems, fundamental to the future of transportation, heavily rely on advanced perception, decision-making, and control technologies. These systems employ a range of sensors [1] such as camera, LiDAR and radar as depicted in Fig. 1, to effectively perceive the surrounding environment. This capability is crucial for recognizing road signs, detecting and tracking vehicles, and predicting pedestrian behavior, enabling safe operation amidst complex traffic conditions.
@@
The primary task of perception is to accurately understand the surrounding environment and minimize collision risks. This is where 3D object detection methods become essential. These approaches enable the autonomous systems to accurately identify objects in the vicinity, including their position, shape, and category [2]. Such detailed environmental perception enhances the system’s ability to comprehend the driving context and make more informed decisions.
@@
The advancement of autonomous driving technologies has spared a wave of research in 3D object detection, leading to the development of diverse and innovative methods. These approaches are typically categorized based on their input types, including Camera-based [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 53, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 17, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100], Point Cloud-based [101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220], and multimodal methods [221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 136, 232, 233, 234, 107, 235, 236, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251]. The current landscape of 3D object detection methods is prolific, necessitating a comprehensive summarization to offer intriguing insights for the research community. While comprehensive, prior surveys [2, 252] often overlook the safety aspects of autonomous driving perception, particularly in terms of the system robustness against varying testing data after deployment.
@@
In real-world testing scenarios, the conditions encountered can greatly differ from those during training, The environmental variability, sensor discrepancies or noise, and spatial misalignment can cause a shift in the input sensory data distribution, leading to a significant drop in detector performance [233, 230, 253, 254]. We identify and discuss three major factors critical for assessing the detection robustness: 1) Environmental Variability: The detection algorithm needs to perform well under different environmental conditions, including variations in lighting, weather, and seasonal changes. The algorithm should exhibit adaptability, ensuring that it does not fail due to changes in the environment. 2) Sensor Noise: This includes handling noise introduced by sensor malfunctions such as motion blur to camera. The algorithm must possess the capability to effectively manage hardware noise, ensuring accurate processing of input data. 3) Misalignment: In real-world scenarios, sensor calibration errors can complicate the synchronization of multimodal input data, causing misalignment due to external factors (e.g., uneven road surfaces) or internal factors (e.g., system clock misalignment). The algorithm should be fault-tolerant and may incorporate an elastic alignment to mitigate misalignment’s impact on detection performance.
@@
To ensure safe operation in varying test environments, assessing the robustness of 3D object detection algorithms is essential. They must maintain efficient, accurate, and reliable performance across diverse scenarios. In this survey, we conduct extensive experimental comparisons among existing algorithms. Centered around ’Accuracy, Latency, Robustness’, we delve into existing solutions, offering insightful guidance for practical deployment in autonomous driving.
@@
Accuracy Current research often prioritizes accuracy as a key performance metric. However, a deeper understanding of these methods’ performance in complex environments and extreme weather conditions is needed to ensure real-world reliability. A more detailed analysis of false positives and false negatives is necessary for improvement.
@@
Latency Real-time capability is vital for autonomous driving. The latency of a 3D object detection method impacts the system’s ability to make timely decisions, particularly in emergencies.
@@
Robustness Robustness refers to the system’s stability under various conditions, including weather, lighting, sensory and alignment changes. Many existing evaluations may not fully consider the diversity of real-world scenarios, necessitating a more comprehensive adaptability assessment.
@@
Through an in-depth analysis of extensive experimental results, with a focus on ‘Accuracy, Latency, Robustness,’ we have identified significant advantages in safety perception with multimodal 3D detection in safety perception. By integrating information from diverse sensors or data sources, the multimodal methods provide a richer and more diverse perception capability for autonomous driving systems, thereby enhancing the understanding and response to the surrounding environment. Our research provides practical guidance for the future deployment of autonomous driving technology. By discussing these key areas, we aim to align the technology more closely with real-world needs and enhance its societal benefits effectively.
@@
The structure of this paper is organized as follows. First, we introduce the datasets and evaluation metrics for 3D object detection, with a particular focus on robustness in Section II. The subsequent sections systematically examine existing 3D object detection methods, including Camera-only (Section III), LiDAR-only (Section IV), and multimodal approaches (Section V). The paper concludes with a comprehensive summary of our findings VI.
@@ II Datasets
+ The dense data format incorporates additional color and texture information. |
- Missing depth information the camera will be affected by light, weather, etc. |
Point cloud |
LiDAR |
@@
+ With accurate depth information less affected by light +larger field of view |
-High computational cost for sparse and disordered point cloud data and no color information. |
Multimodal |
Camera, LiDAR |
+ Simultaneous color and depth information |
- Fusion methods can produce noise interference |
| Dataset | Year | Sensors | Data Size | Diversity | Scenes | ||
| Frame | Annotation | Scenes | Category | ||||
| KITTI[255] | 2012 | Camera, LiDAR | 15K | 200K | 50 | 3 | Daytime and sunny days only |
| nuScenes[256] | 2019 | Camera, LiDAR, Radar | 40K | 1.4M | 1000 | 10 | Night and rainy days |
| Lyft L5[257] | 2019 | Camera, LiDAR | 46K | 1.3M | 366 | 9 | Daytime and sunny days only |
| H3D[258] | 2019 | LiDAR | 27K | 1.1M | 160 | 8 | Daytime and sunny days only |
| Appllo[259] | 2019 | Camera, LiDAR | 140K | - | 103 | 27 | Night and rainy days |
| Argoverse[260] | 2019 | Camera, LiDAR | 46K | 993K | 366 | 9 | Night and rainy days |
| A*3D [261] | 2019 | Camera, LiDAR | 39K | 230K | - | 7 | Night and severely obscured data |
| Waymo[262] | 2020 | Camera, LiDAR | 230K | 12M | 1150 | 3 | Night and rainy days |
| A2D2 [263] | 2020 | Camera, LiDAR | 12.5K | 43K | - | 38 | Daytime and sunny days only |
| PandaSet [264] | 2020 | Camera, LiDAR | 14K | - | 179 | 28 | Daytime and sunny days only |
| KITTI-360 [265] | 2020 | Camera, LiDAR | 80K | 68K | 11 | 19 | Daytime and sunny days only |
| Cirrus [266] | 2020 | Camera, LiDAR | 6285 | - | 12 | 8 | Daytime and sunny days only |
| ONCE [267] | 2021 | Camera, LiDAR | 15K | 417K | - | 5 | Night and rainy days |
| OpenLane [268] | 2022 | Camera, LiDAR | 200K | - | 1000 | 14 | Daytime and sunny days only |
@@
Currently, autonomous driving systems primarily rely on sensors such as cameras, LiDAR, and radar, generating data in two modalities: point clouds and images. Based on these data types, existing public benchmarks predominantly manifest in three forms: Camera-only, LiDAR-only, and multimodal. Table I delineates the advantages and disadvantages of each of these three forms. Among them, there are many reviews [269, 252, 270, 271, 272, 273, 274, 275] providing a comprehensive overview of clean autonomous driving datasets as shown in II, The most notable ones include KITTI[255], nuScenes[256], and Waymo[262].
@@
In recent times, the pioneering work on clean autonomous driving datasets has provided rich resources for 3D object detection. As autonomous driving technology transitions from breakthrough stages to practical implementation, we have undertaken some guided research to review the currently available robustness datasets systematically. We focus more on noisy scenarios and have systematically reviewed datasets related to the robustness of 3D detection. Many studies collect new datasets to evaluate model robustness under different conditions. Early research has explored camera-only approaches under adverse conditions [276, 277], with datasets notably small in scale and exclusively applicable to camera-only visual tasks rather than multimodal sensor stacks that include LiDAR. Subsequently, a series of multimodal datasets [278, 279, 280, 281] focus on noise concerns. For instance, the GROUNDED dataset [278] focuses on ground-penetrating radar localization under varying weather conditions. Additionally, the ApolloScape open dataset [280] incorporates LiDAR, camera, and GPS data, encompassing cloudy and rainy conditions and brightly lit scenarios.
@@
Due to the prohibitive cost of collecting extensive noisy datasets from the real world, rendering the formation of large-scale datasets impractical, many studies have shifted their focus to synthetic datasets. ImageNet-C [282] is a seminal work in corruption robustness research, benchmarking classical image classification models against prevalent corruptions and perturbations. This line of research has subsequently extended to include robustness datasets tailored for 3D object detection in autonomous driving. Additionally, there are adversarial attacks [283, 284, 285] designed for studying the robustness of 3D object detection. However, these attacks may not exclusively concentrate on natural corruption, which is less prevalent in autonomous driving scenarios. To better emulate the distribution of noise data in the real world, several studies [254, 286, 287, 253, 288, 289] have developed toolkits for robustness benchmarks. These benchmark toolkits [254, 286, 287, 253, 288, 289] enable the simulation of various scenarios using clean autonomous driving datasets, such as KITTI [255], nuScenes [256], and Waymo [262]. Among them, Dong et al.[254] systematically designed 27 common corruptions in 3D object detection to benchmark the corruption robustness of existing detectors. By applying these corruptions comprehensively on public datasets, they established three corruption-robust benchmarks: KITTI-C, nuScenes-C, and Waymo-C. [254] denote model performance on the original validation set as . For each corruption type at each severity , [254] adopt the same metric to measure model performance as . The corruption robustness of a model is calculated by averaging over all corruption types and severities as
@@
Where is the set of corruptions in evaluation, note that for different kinds of 3D object detectors, the set of corruptions can be different (e.g., [254] do not evaluate camera noises for LiDAR-only models). Thus, the results of APcor are not directly comparable between different kinds of models, and [254] performs a fine-grained analysis under each corruption. [254] also calculates relative corruption error (RCE) by measuring the percentage of performance drop as
@@
Unlike KITTI-C and Waymo-C, nuScenes-C primarily assesses performance using the mean Average Precision (mAP) and nuScenes Detection Score (NDS) computed across ten object categories. The mAP is determined using the 2D center distance on the ground plane instead of the 3D Intersection over Union (IoU). The NDS metric consolidates mAP with other aspects, such as scale and orientation, into a unified score. Analogous to KITTI-C, Ref. [254] denote the model’s performance on the validation set as and , respectively. The corruption robustness metrics, and , are evaluated by averaging over all corruption types and severities. Additionally, Ref. [254] calculates the Relative Corruption Error (RCE) under both mAP and NDS metrics, similar to the formulation in Eq.2.
@@
Additionally, some studies [283, 286, 290] examine robustness in single-modal contexts. For instance, Ref. [286] proposes a LiDAR-only benchmark that utilizes physical-aware simulation methods to simulate degraded point clouds under various real-world common corruptions. This benchmark, tailored for point cloud detectors, includes 1,122,150 examples across 7,481 scenes, covering 25 common corruption types with six severity levels. Moreover, Ref. [286] devise a novel evaluation metrics, including and mCE. Ref. [286] calculates corruption error (CE) to assess performance degradation based on Overall Accuracy (OA) by:
@@
where is the overall accuracy of detector m under corruption of severity level (exclude ”clean,” i.e., severity level 0) and clean represent the clean data. For detection m, we can calculate the mean CE () for each detector by:
III Camera-based 3D Object Detection

### v2

@@ Abstract
In the realm of modern autonomous driving, the perception system is indispensable for accurately assessing the state of the surrounding environment, thereby enabling informed prediction and planning. The key step to this system is related to 3D object detection that utilizes vehicle-mounted sensors such as LiDAR and cameras to identify the size, the category, and the location of nearby objects. Despite the surge in 3D object detection methods aimed at enhancing detection precision and efficiency, there is a gap in the literature that systematically examines their resilience against environmental variations, noise, and weather changes. This study emphasizes the importance of robustness, alongside accuracy and latency, in evaluating perception systems under practical scenarios. Our work presents an extensive survey of camera-only, LiDAR-only, and multi-modal 3D object detection algorithms, thoroughly evaluating their trade-off between accuracy, latency, and robustness, particularly on datasets like KITTI-C and nuScenes-C to ensure fair comparisons. Among these, multi-modal 3D detection approaches exhibit superior robustness, and a novel taxonomy is introduced to reorganize the literature for enhanced clarity. This survey aims to offer a more practical perspective on the current capabilities and the constraints of 3D object detection algorithms in real-world applications, thus steering future research towards robustness-centric advancements.
@@
3D Object Detection, Perception, Robustness, Autonomous DrivingI Introduction
@@
AUTONOMOUS driving systems, fundamental to the future of transportation, heavily rely on advanced perception, decision-making, and control technologies. These systems employ a range of sensors [1] such as camera, LiDAR and radar as depicted in Fig. 1, to effectively perceive surrounding environments. This capability is crucial for recognizing road signs, detecting and tracking vehicles, and predicting pedestrian behaviors, enabling safe operations amidst complex traffic conditions [2].
@@
The primary task of perception is to accurately understand the surrounding environment and minimize collision risks [3]. This is where 3D object detection methods become essential. These approaches enable the autonomous systems to accurately identify objects in the vicinity, including their position, shape, and category [4]. Such detailed environmental perception enhances the system’s ability to comprehend the driving context and make more informed decisions.
@@
The advancement of autonomous driving technologies has spared a wave of research in 3D object detection, leading to the development of diverse and innovative methods. These approaches are typically categorized based on their input types, including camera-only [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], LiDAR-only [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98], and multi-modal methods [99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 65, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122]. The current landscape of 3D object detection methods is prolific, necessitating a comprehensive summarization to offer intriguing insights to the research community. While comprehensive, prior surveys, such as [4, 123], often overlook the safety aspects of autonomous driving perception, particularly in terms of the system’s robustness against varying testing data following deployment.
@@
In real-world testing scenarios, the conditions encountered usually greatly differ from those during training. The environmental variability, sensor discrepancies or noise, and spatial misalignment can cause a shift in the input sensory data distribution, leading to a significant drop in detector performance [111, 108, 124, 125]. We identify and discuss three major factors critical for assessing detection robustness.
Environmental Variability. A detection algorithm needs to perform well under different environmental conditions, including variations in lighting, weather, and seasons. The algorithms should exhibit adaptability, ensuring that it does not fail due to changes in the environment.
Sensor Noise. This includes handling noise introduced by sensor malfunctions, such as motion blur in a camera. An algorithm must possess the capability to effectively manage hardware noise, ensuring the accurate processing of input data.
Misalignment. In real-world scenarios, sensor calibration errors can complicate the synchronization of multi-modal input data, causing misalignment due to external factors (e.g., uneven road surfaces) or internal factors (e.g., system clock misalignment). An algorithm should be fault-tolerant and may incorporate an elastic alignment to mitigate the impact of misalignment on detection performance.
@@
To ensure safe operation in varying test environments, assessing the robustness of 3D object detection algorithms is essential. They must maintain efficient, accurate, and reliable performance across diverse scenarios. In this survey, we conduct extensive experimental comparisons among existing algorithms. Centered around ‘Accuracy, Latency, Robustness’, we delve into existing solutions, offering insightful guidance for practical deployment in autonomous driving.
@@
Accuracy: Current researches often prioritize accuracy as a key performance metric. However, a deeper understanding of these methods’ performance in complex environments and extreme weather conditions is needed to ensure real-world reliability. A more detailed analysis of false positives and false negatives is necessary for improvement.
@@
Latency: Real-time capability is vital for autonomous driving. The latency of a 3D object detection method impacts the system’s ability to make timely decisions, particularly in emergencies.
@@
Robustness: Robustness refers to the system’s stability under various conditions, including weather, lighting, sensory and alignment changes. Many existing evaluations may not fully consider the diversity of real-world scenarios, necessitating a more comprehensive adaptability assessment.
@@
Through an in-depth analysis of extensive experimental results, with a focus on ‘Accuracy, Latency, Robustness’, we have identified significant advantages in safety perception with Multi-modal 3D detection in safety perception. By integrating information from diverse sensors or data sources, Multi-modal methods provide a richer and more diverse perception capability for autonomous driving systems, thereby enhancing the understanding and responding to the surrounding environment. Our research provides practical guidance for the future deployment of autonomous driving technology. By discussing these key areas, we aim to align the technology more closely with real-world needs and enhance its societal benefits effectively.
@@
The structure of this paper is organized as follows: First, we introduce the datasets and evaluation metrics for 3D object detection, with a particular focus on robustness in Section II. Subsequent sections systematically examine existing 3D object detection methods, including camera-only approaches (Section III), LiDAR-only approaches (Section IV), and multi-modal approaches (Section V). The paper concludes with a comprehensive summary of our findings in Section VII.
@@ II Datasets
| Image | Camera | ~ | + The dense data format incorporates additional color and texture information. | - Missing depth information the camera will be affected by light, weather, etc. |
@@
| Point cloud | LiDAR | ~ | + With accurate depth information less affected by light +larger field of view | -High computational cost for sparse and disordered point cloud data and no color information. |
| Multi-modal | Camera, LiDAR | ~ | + Simultaneous color and depth information | - Fusion methods can produce noise interference |
@@
Currently, autonomous driving systems primarily rely on sensors such as cameras, and LiDAR, generating data in two modalities, point clouds and images. Based on these data types, existing public benchmarks predominantly manifest in three forms: camera-only, LiDAR-only, and multi-modal. Table I delineates the advantages and the disadvantages of each of these three forms. Among them, there are many reviews [126, 123, 127, 128, 129, 130, 131, 132] providing a comprehensive overview of clean autonomous driving datasets as shown in Table II. The most notable ones include KITTI[133], nuScenes[134], and Waymo[135].
@@
In recent times, the pioneering work on clean autonomous driving datasets has provided rich resources for 3D object detection. As autonomous driving technology transitions from breakthrough stages to practical implementation, we have conducted some guided researches to systematically review the currently available robustness datasets. We have focused more on noisy scenarios and systematically reviewed datasets related to the robustness of 3D detection. Many studies have collected new datasets to evaluate model robustness under different conditions. Early research explored camera-only approaches under adverse conditions [136, 137], with datasets that were notably small in scale and exclusively applicable to camera-only visual tasks, rather than multi-modal sensor stacks that include LiDAR. Subsequently, a series of multi-modal datasets [138, 139, 140, 141] have focused on noise concerns. For instance, the GROUNDED dataset [138] focuses on ground-penetrating radar localization under varying weather conditions. Additionally, the ApolloScape open dataset [140] incorporates LiDAR, camera, and GPS data, encompassing cloudy and rainy conditions as well as brightly lit scenarios. The Ithaca365 dataset [141] is designed for robustness in autonomous driving research, providing scenarios under various challenging weather conditions, such as rain and snow.
@@
Due to the prohibitive cost of collecting extensive noisy datasets from the real world, rendering the formation of large-scale datasets impractical, many studies have shifted their focus to synthetic datasets. ImageNet-C [142] is a seminal work in corruption robustness research, benchmarking classical image classification models against prevalent corruptions and perturbations. This line of research has subsequently been extended to include robustness datasets tailored for 3D object detection in autonomous driving. Additionally, there are adversarial attacks [143, 144, 145] designed for studying the robustness of 3D object detection. However, these attacks may not exclusively concentrate on natural corruption, which is less common in autonomous driving scenarios.
To better emulate the distribution of noise data in the real world, several studies [125, 146, 147, 124, 148, 149] have developed toolkits for robustness benchmarks. These benchmark toolkits [125, 146, 147, 124, 148, 149] enable the simulation of various scenarios using clean autonomous driving datasets, such as KITTI [133], nuScenes [134], and Waymo [135]. Among them, Dong et al.[125] systematically designed 27 common corruptions in 3D object detection to benchmark the corruption robustness of existing detectors. By applying these corruptions comprehensively on public datasets, they established three corruption-robust benchmarks: KITTI-C, nuScenes-C, and Waymo-C. [125] denotes model performance on the original validation set as . For each corruption type at each severity , [125] adopts the same metric to measure model performance as . The corruption robustness of a model is calculated by averaging over all corruption types and severities as
@@
Where is the set of corruptions in evaluation. It should be noticed that for different kinds of 3D object detectors, the set of corruptions can be different (e.g., [125] has not evaluated camera noises for LiDAR-only models). Thus, the results of APcor are not directly comparable between different kinds of models, and [125] performs a fine-grained analysis under each corruption. It also calculates relative corruption error (RCE) by measuring the percentage of performance drop as
@@
Unlike KITTI-C and Waymo-C, nuScenes-C primarily assesses performance using the mean Average Precision (mAP) and nuScenes Detection Score (NDS) computed across ten object categories. The mAP is determined using the 2D center distance on the ground plane instead of the 3D Intersection over Union (IoU). The NDS metric consolidates mAP with other aspects, such as scale and orientation, into a unified score. Analogous to KITTI-C, [125] denotes the model’s performance on the validation set as and , respectively. The corruption robustness metrics, and , are evaluated by averaging over all corruption types and severities. Additionally, [125] calculates the Relative Corruption Error (RCE) under both mAP and NDS metrics, similar to the formulation in Eq.2.
| Dataset | Year | Sensors | Data Size | Diversity | ||
| Frame | Annotation | Scenes | Category | |||
| KITTI[133] | 2012 | CL | 15K | 200K | 50 | 3 |
| nuScenes[134] | 2019 | CLR | 40K | 1.4M | 1000 | 10 |
| Lyft L5[150] | 2019 | CL | 46K | 1.3M | 366 | 9 |
| H3D[151] | 2019 | L | 27K | 1.1M | 160 | 8 |
| Appllo[152] | 2019 | CL | 140K | - | 103 | 27 |
| Argoverse[153] | 2019 | CL | 46K | 993K | 366 | 9 |
| A*3D [154] | 2019 | CL | 39K | 230K | - | 7 |
| Waymo[135] | 2020 | CL | 230K | 12M | 1150 | 3 |
| A2D2 [155] | 2020 | CL | 12.5K | 43K | - | 38 |
| PandaSet [156] | 2020 | CL | 14K | - | 179 | 28 |
| KITTI-360 [157] | 2020 | CL | 80K | 68K | 11 | 19 |
| Cirrus [158] | 2020 | CL | 6285 | - | 12 | 8 |
| ONCE [159] | 2021 | CL | 15K | 417K | - | 5 |
| OpenLane [160] | 2022 | CL | 200K | - | 1000 | 14 |
@@
Additionally, some studies [143, 146, 161] examine robustness in single-modal contexts. For instance, [146] proposes a LiDAR-only benchmark that utilizes physically-aware simulation methods to simulate degraded point clouds under various real-world common corruptions. This benchmark, tailored for point cloud detectors, includes 1,122,150 examples across 7,481 scenes, covering 25 common corruption types with six severity levels. Moreover, [146] devises a novel evaluation metric, including and mCE, and calculates corruption error (CE) to assess performance degradation based on Overall Accuracy (OA) by
@@
where is the overall accuracy of detector m under corruption of severity level (excluding “clean,” i.e., severity level 0) and clean represents the clean data. For detector m, we can calculate the mean CE () for each detector by
III Camera-only 3D Object Detection